Chemcial Equation: C3H8 + O2 ( CO2 + H2O




If 50.0 grams of Cl2 gas is reacted with enough sodium hydroxide (NaOH), calculate the following:




Mass in grams of sodium hypochlorite (NaOCL) produced




Moles of sodium chloride (NaCl) produced




Chemical Equation: NaOH + Cl2 ( NaOCl + NaCl + H2O

1
Expert's answer
2022-04-25T17:02:03-0400

2NaOH + Cl2 = NaOCl + NaCl + H2O

n(Cl2) = m(Cl2) / Mr(Cl2) = 50.0 g / 71.0 g/mol = 0.704 mol

1 mol Cl2 - 1 mol NaOCl

0.704 mol Cl2 - x mol NaOCl

x = 0.704 mol

m(NaOCl) = n(NaOCl) * Mr(NaOCl) = 0.704 mol * 74.5 g/mol = 52.45 g

1 mol Cl2 - 1 mol NaCl

0.704 mol Cl2 - x mol NaCl

x = 0.704 mol